Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
cifaldi, nardecchia - 20:42 Friday 29 August 2025 (67596) Print this report
NI DAS outer ring power step reverted

This morning, Maria was contacted by the control room and asked to revert the change in power of the NI DAS outer ring that had been performed on Tuesday (67571). She carried out the change at around 9:00 UTC. 

I’m not sure if this step improved the situation, since the operator’s entry also mentions bad weather (67592).

The power on the pickoff is now 0.502 W and 3.1 W injected  into the ITF.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×